OpenStudy (queenof17):

The midpoint of A (-4, 2) and B(8, 5) is (2, 7/2) (-6, 3/2) (4, 7/2)

OpenStudy (freckles):

just find the average of -4 and 8 and just find the average of 2 and 5

OpenStudy (freckles):

the average of two numbers let's call the numbers a and b is (a+b)/2

OpenStudy (freckles):

so you sum the numbers and divide the result by 2

OpenStudy (queenof17):

so it's 2, 7/2

OpenStudy (queenof17):

?

OpenStudy (freckles):

yes the average of -4 and 8 is (-4+8)/2 =4/2 =2 and the average of 2 and 5 is (2+5)/2 =7/2 so the midpoint is (2,7/2)

OpenStudy (queenof17):

okay, thank you very, very much!!